OpenStudy (anonymous):

What is the domain of the function f(x) = x+1/2x-1?

OpenStudy (campbell_st):

well there is a restriction in the domain, caused by the denominator the domain is all real x except where 2x - 1 = 0 so you need to calculate the restriction... remember, zero denominators are undefined..

OpenStudy (anonymous):

so it'd be all real numbers except -1?

OpenStudy (campbell_st):

nope... what value does x take if 2x - 1 = 0 and its not -1 just solve the equation for x

OpenStudy (anonymous):

x = 1/2?

OpenStudy (campbell_st):

that's the restriction